OpenStudy (anonymous):

A meter stick is held perpendicular to the ground. it casts a shadow 1.5 m long. At teh same time, a telephone pole casts a shadow that is 9m long. How tall is the telephone pole?

OpenStudy (anonymous):

13.5 meters

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

well if look at the meter stick... a meter is is equal to 1 meter, and it's shadow is 1.5 meters... so the ratio is 1:1.5 ^ you can simplify the ratio to: 2:3 ^ just divide everything by .5 now if 2:3 equals 9:x what would x be? Well.. find half of 9.. which is 4.5 and add that to 9 and you get 13.5.

OpenStudy (anonymous):

thanks!!! :)

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

\[1/1.5=x/9\rightarrow x=6\]

OpenStudy (anonymous):

thanks AA! :)

OpenStudy (anonymous):

madscientist had the right idea, but set up one side of the proportion inverted.